OpenStudy (anonymous):

Prove that the only limit points of A are {0,1} and the isolated points of A are A\{0,1}

OpenStudy (anonymous):

\[A = \left\{ \frac{ n+1 }{ 2n } +\frac{ (-1)^{n} }{ 2 }, n \in \mathbb{N} \right\} \] A continuation from my previous question. I know the limit and isolated points, just need help showing it is all :)